# The trade-off between diagonal and off-diagonal elements in the eigenstate thermalization hypothesis

### v2

@@ II.2 Weak ETH with eigenstate typicality
Thus, when $\Delta(\rho,A)$ is very small, the expectation of the projectively measured observable would concentrate on the expectation of observable calculated with respect to the state $\rho_{B}$. If
$\Delta(\rho^{\text{mc}},A_{B_{1}})\sim\mathcal{O}(N^{-\alpha}),$ (9)
with $10<\alpha<$ 1, meaning that the expectations of a local observable $A^{B_{1}}$ with respect to the results of local measurements concentrate around the expectation of $A^{B_{1}}$ with respect to $\rho^{\text{mc}}$, then it follows that each pure state $\Pi^{i}_{B}$ cannot be distinguished from the microcanonical state $\rho^{\text{mc}}$ in the large $N$ limit. This result characterizes the weak ETH with eigenstate typicality [6, 11]. By substituting the measure $d^{(1)}$ with other measures, different forms of the weak ETH can be obtained.
II.3 The average observable
In a many-body system, the states of individual subsystems exhibit fluctuations. Even in a system with a Hamiltonian that is translationally invariant, only certain states, such as the Gibbs state (canonical ensemble) and the microcanonical ensemble, maintain guaranteed translational invariance. The energy eigenstates themselves can still fluctuate. Consequently, compared to the translationally invariant ensemble, spatial fluctuations can make the energy eigenstates more distinguishable.
Consider partitioning the lattice of system $B$ into $\mathcal{C}=\lvert B\rvert/\lvert B_{1}\rvert$ blocks of equal size, where $\lvert B\rvert$ denotes the number of lattice points in $B$. These $\mathcal{C}$ blocks are identical copies of $B_{1}$. In order to eliminate the influence of spatial fluctuations, or consider the macroscopic observable [12], one can use the average observable
$A^{B}=\frac{1}{\mathcal{C}}\sum_{i}A^{B_{i}},$ (10)
where $A^{B_{i}}$ represents the translational copy of $A^{B_{1}}$. Substituting $A^{B}$ into eq. 1 gives
@@ II.3 The average observable
where $\sigma^{j,k}_{B_{1}}$ are the translational copies of $\sigma^{j}_{B_{k}}$, $\overline{\rho}_{B}=(\sum_{k}\hat{T}_{k}\rho_{B})/\mathcal{C}$ and $\hat{T}_{k}$ is the translation operator that maps $B_{k}$ to $B_{1}$. According to our previous work [16], an upper bound can be given by the variance
$d^{(1)}(\Pi^{ij}_{B},\rho;A)^{2}\leqslant V(\overline{\rho},\overline{O}^{B_{1}}_{ij})\lVert A^{B_{1}}\rVert_{\infty}^{2},$ (12)
where $\overline{O}^{B_{1}}_{ij}=\frac{1}{\mathcal{C}}\sum_{k}\mathcal{J}_{\rho_{B_{1}}}^{-1/2}(\sigma^{ij,k}_{B_{1}})$, and $\sigma^{ij,k}_{B_{1}}$ are the translational copies of $\sigma^{ij}_{B_{k}}$.
@@ III.1 Trade-off between diagonal and off-diagonal elements
The diagonal element $V_{\text{dg}}^{i}=V(\rho_{B},O^{B_{1}}_{ii})$ and the off-diagonal element $V_{\text{off}}^{ij}=V(\rho_{B},O^{B_{1}}_{ij;i\neq j})$ in $d^{(3)}$ are not independent. According to appendix A, they are related by the following equation
$V_{\text{dg}}^{i}+\sum_{j;j\neq i}V_{\text{off}}^{ij}=\text{Tr}(\rho_{B_{1}}^{-1})-1,$ (13)
where the index $j$ sums over the entire Hilbert space except $i$, and its dimension $D_{N}-1$ grows exponentially with the number of sites. In contrast, the right-hand side of (13) depends solely on the local state of $\rho_{B}$, generally increasing with $D_{B_{1}}^{2}$ but remaining relatively constant with $N$. This causes the off-diagonal elements to be exponentially suppressed. Considering the average of these off-diagonal elements
$V_{\text{off}}^{i}=\frac{1}{D_{N}-1}\sum_{j;j\neq i}V_{\text{off}}^{ij},$ (14)
which will decrease exponentially with $N$. However, this exponential decrease in the average of the off-diagonal elements does not imply that the maximum values of the off-diagonal elements and the diagonal element will also decrease exponentially. Although the diagonal elements are not directly suppressed, eq. 13 provides a trade-off relation between the diagonal element and the off-diagonal elements. If $V_{\text{dg}}^{i}/V_{\text{off}}^{i}=f^{i}(N)$, then we have
@@
On one hand, if $f^{i}(N)$ does not increase exponentially, or grows exponentially at a rate much slower than $D_{N}$, then in the large $N$ limit, the diagonal elements will also be exponentially small. On the other hand, if the diagonal element is not decreasing exponentially, $f^{i}(N)$ must be growing exponentially. Thus, the exponential growth of $f^{i}(N)$ is a necessary condition for the diagonal element to avoid exponential decay.
@@
$f^{i}(N)\leq\frac{D_{N}-1}{\sum_{j;j\neq i}\lVert(\sigma^{i}_{B_{1}}-\rho_{B_{1}})(\sigma^{ij}_{B_{1}})^{-1}\rVert^{-2}_{\infty}}.$ (17)
@@
This inequality’s right-hand side may provide a more straightforward framework to discuss the effects of interactions and the choice of the $\rho_{B_{1}}$ on $f^{i}(N)$.
@@
Figure 1: Diagonal and off-diagonal elements versus eigenstate. The eigenstate indices are arranged from low to high energy. In the plot, $\text{Var}_{\text{dg}}=V_{\text{dg}}^{i}$, $\text{Var}_{\text{off}}=\sum_{j;j\neq i}V_{\text{off}}^{ij}$, $\text{SN}_{\text{dg}}=\lVert\sigma^{B_{1}}_{i}-\rho_{B_{1}}\rVert_{1}^{2}$ and $\text{SN}_{\text{off}}=\sum_{j;j\neq i}\lVert\sigma^{B_{1}}_{ij}\rVert_{1}^{2}$. The measure $\text{Var}_{\text{dg}}$ closely aligns with $\text{SN}_{\text{dg}}$, making the data points for $\text{Var}_{\text{dg}}$ overlap with those for $\text{SN}_{\text{dg}}$. Parameters: $g=1.05$, $h=0.1$, $N=12$, $N_{B_{1}}$=1, $\rho_{B}=\sum_{i}\Pi_{i}/D_{N}$.
@@
The index $i$ in eq. 13 can be chosen arbitrarily, but the right side of the formula is independent of $i$. When the measure $V_{\text{dg}}$ of a given eigenstate is large, the average measure of its off-diagonal elements $V_{\text{off}}^{i}$ must be small, and vice versa. This trade-off is illustrated in fig. 1. Since the measure $d^{(2)}$ and $d^{(3)}$ are related by section II.1, the trade-off between the diagonal and off-diagonal elements is also reflected in $d^{(2)}$, although it is not an exact equation relationship like eq. 13. Calculations using a specific model are shown in fig. 1, where the diagonal element and the sum of the corresponding off-diagonal elements are compared. From the fig. 1, we observe that the diagonal element is generally much larger than the average of the off-diagonal elements. However, because there are $D_{N}-1$ off-diagonal elements, their total sum significantly exceeds that of the diagonal elements.
@@
The formula (13) holds for any $\rho_{B}$ as long as $\rho_{B_{1}}$ is strictly positive. This includes not only the canonical ensemble $\rho^{\text{c}}=e^{-\beta H}/Z(\beta)$ and microcanonical ensembles, but also pure states such as $\Pi_{i}^{B}=\ket{E_{i}}\bra{E_{i}}$ |. In the case of pure states, the diagonal element in eq. 13 will be zero, leading to
$\sum_{j;j\neq i}V(\Pi_{i}^{B},O^{B_{1}}_{ij})=\text{Tr}(\sigma_{B_{1}}^{-1})-1,$ (18)
@@
where the sum of all corresponding off-diagonal elements is directly determined by the local state $\sigma_{B_{1}}=\text{Tr}_{\bar{B}_{1}}\Pi_{i}^{B}$.
@@ III.2 Weak ETH
We examine the use of measure $d^{(3)}$ to study the weak ETH with eigenstate typicality by considering both ”diagonal” and ”off-diagonal” probabilistic typicalities. The ”diagonal” probabilistic typicality is given by
@@
while the ”off-diagonal” probabilistic typicality is
@@
As discussed previously, $\braket{V_{\text{off}}}$ decreases exponentially with $N$. The behavior of $\braket{V_{\text{dg}}}$ depends on the specific model.
@@ III.3 The average observable
If we consider the distinguishability measure $V(\overline{\rho}_{B},\overline{O}^{B_{1}}_{ij})$ given by the average observable, then according to appendix A, the corresponding diagonal and off-diagonal elements exhibit the following trade-off relationship
$V(\rho_{B},\overline{O}^{B_{1}}_{ii})+\sum_{j;j\neq i}V(\rho_{B},\overline{O}^{B_{1}}_{ij})=\frac{1}{\mathcal{C}}(\text{Tr}(\rho_{B_{1}}^{-1})-1)$
$+\frac{1}{\mathcal{C}^{2}}\sum_{\beta,\alpha,l,k;k\neq l}p^{\prime}_{\alpha}\text{Tr}[O^{B_{k},{\dagger}}_{\alpha\beta,1}\otimes O^{B_{l}}_{\alpha\beta,1}(\sigma^{i}_{B_{k}B_{l}}-\sigma^{i}_{B_{k}}\otimes\rho^{1}_{B_{l}})],$ (22)
where $\rho^{1}_{B_{l}}$ is the translational copy of $\rho_{B_{1}}$. Although formula (III.3) is valid for any $\rho_{B}$, it connects to eq. 12 specifically when using $\overline{\rho}_{B}$. Thus, all quantities related to $\rho_{B}$ should be replaced with $\overline{\rho}_{B}$, which is translation invariant. In this case $\overline{\rho}^{1}_{B_{l}}=\overline{\rho}_{B_{l}}$. The local term in section III.3 becomes $\text{Tr}(\overline{\rho}_{B_{1}}^{-1})-1$, which is related to the local state $\overline{\rho}_{B_{1}}$ and includes an additional factor of $1/\mathcal{C}$. The local term decreases algebraically as $\mathcal{C}$ increases. The second term on the right side of section III.3 is related to the spatial correlation of $\sigma^{i}$ and the difference between $\sigma^{i}_{B_{l}}$ and $\overline{\rho}_{B_{l}}$.
@@
$\braket{V_{\text{dg}}}=\sum_{i}p_{i}V(\rho,\overline{O}^{B_{1}}_{ii}).$ (23)
@@
$\braket{V_{\text{off}}}=\frac{1}{D_{N}-1}\sum_{i,j;j\neq i}p_{i}V(\rho,\overline{O}^{B_{1}}_{ij}).$ (24)
@@
$V^{\text{Avg}}_{\text{Cor}}:=\frac{1}{\mathcal{C}^{2}}\sum_{\beta,\alpha,l,k;k\neq l}p^{\prime}_{\alpha}\text{Tr}[O^{B_{k},{\dagger}}_{\alpha\beta}\otimes O^{B_{l}}_{\alpha\beta}(\rho_{B_{k}B_{l}}-\rho_{B_{k}}\otimes\rho_{B_{l}})].$ (26)
In formula (24), except for $\rho_{B_{k}B_{l}}$ and $\rho_{B_{k}}$, all quantities related to $\rho_{B}$ need to be replaced with $\overline{\rho}_{B}$. If $\rho_{B}$ is translation invariant, then $\overline{\rho}_{B}=\rho_{B}$, and the second term on the right side of eq. 25 is solely related to the correlation of $\rho_{B}$. If the spatial correlation of $\rho_{B}$ decays exponentially or algebraically fast enough, the right side of eq. 25 will exhibit algebraic decay of the form $O(N^{-\epsilon})$ [16]. Compared to eq. 21, the right side of eq. 25 depends not only on the local state but also on the rate of correlation decay. Both terms generally decrease algebraically with increasing $N$, consistent with the intuition that spatial averaging smooths and reduces differences. Although this algebraic decay alone does not guarantee strong ETH, it supports the validity of weak ETH even for integrable systems. This algebraic decay aligns with the findings of our previous article [16], though we did not previously account for the $D_{N}-1$ off-diagonal elements in eq. 25, which cause the average off-diagonal elements to be exponentially suppressed and further restrict their values. Consequently, as the right side of eq. 25 decreases with $N$, the exponential decrease of $\braket{V_{\text{off}}}$ remains applicable. Additionally, due to the smoothing effect of the average observable, the values of $\braket{V_{\text{dg}}}$ and $\braket{V_{\text{off}}}$ here will be smaller than those defined in the previous section.

### v3

@@ II.3 Weak ETH with eigenstate typicality
Thus, when $\Delta(\rho,A)$ is very small, the expectation of the projectively measured observable would concentrate on the expectation of observable calculated with respect to the state $\rho_{B}$. For any microcanonical ensemble $\rho^{\text{mc}}$, if
$\Delta(\rho^{\text{mc}},A^{B_{1}})\sim\mathcal{O}(N^{-\alpha}),$ (18)
with $0<\alpha<1$, which means that the expectations of a local observable $A^{B_{1}}$ with respect to the results of local measurements concentrate around the expectation of $A^{B_{1}}$ with respect to $\rho^{\text{mc}}$, then it follows that each pure state $\Pi^{i}_{B}$ cannot be distinguished from the microcanonical state $\rho^{\text{mc}}$ in the large $N$ limit. This result characterizes the weak ETH with eigenstate typicality [11]. By substituting the measure $d^{(1)}$ with other measures, different forms of the weak ETH can be obtained.
II.4 The average observable
Subsystems in many-body quantum systems inherently exhibit state fluctuations. Even when governed by translationally invariant Hamiltonians, only equilibrium ensembles like the Gibbs state (canonical ensemble) and microcanonical ensemble preserve strict translational symmetry. Individual energy eigenstates, however, manifest residual spatial fluctuations that enhance their distinguishability compared to translation-invariant ensembles.
[…]
$A^{B}:=\frac{1}{\mathcal{C}}\sum_{i}A^{B_{i}},$ (19)
where $A^{B_{i}}$ represents the translational copy of $A^{B_{1}}$. This regularization procedure effectively projects local operators onto the translation-symmetric subspace, reconciling eigenstate fluctuations with thermodynamic predictability. Substituting $A^{B}$ into Eq. 1 gives
@@ II.4 The average observable
where $\sigma^{ij,k}_{B_{1}}$ is the translational copy of $\sigma^{ij}_{B_{k}}$ and $\overline{\rho}_{B_{1}}:=\text{Tr}_{\bar{B}_{1}}\overline{\rho}_{B}$. The density matrix $\overline{\rho}_{B}:=(\sum_{k}\hat{T}_{k}\rho_{B})/\mathcal{C}$ and $\hat{T}_{k}$ is the translation operator that maps $B_{k}$ to $B_{1}$. According to Sections II.4 and II.1, an upper bound can be given by
$[d^{(1)}(\Pi^{ij}_{B},\rho;A^{B})]^{2}$
$\leq\lVert(\frac{1}{\mathcal{C}}\sum_{k}\sigma^{ij,k}_{B_{1}}-\delta_{ij}\overline{\rho}_{B_{1}})\overline{\rho}^{-1/2}_{B_{1}}\overline{\rho}^{1/2}_{B_{1}}(A^{B_{1}}-\braket{A^{B_{1}}}_{\overline{\rho}_{B}})\rVert_{1}^{2}$
$\leqslant V(\overline{\rho}_{B},\overline{O}^{B_{1}}_{ij})\times V(\overline{\rho}_{B},A^{B_{1}}),$ (21)
where $\overline{O}^{B_{1}}_{ij}:=\frac{1}{\mathcal{C}}\sum_{k}\mathcal{J}_{\overline{\rho}_{B_{1}}}^{-1/2}(\sigma^{ij,k}_{B_{1}})$, and $\sigma^{ij,k}_{B_{1}}$ are the translational copies of $\sigma^{ij}_{B_{k}}$.
@@ III Trade-off between diagonal and off-diagonal elements
In this section, we examine the relationship between the diagonal elements $d(\sigma_{B_{1}}^{ii},\rho_{B_{1}})$ and the off-diagonal elements $d(\sigma_{B_{1}}^{ij;i\neq j},\rho_{B_{1}})$ in the context of the ETH metric. For clarity, we emphasize that throughout this work, the terms “diagonal elements” and “off-diagonal elements” specifically refer to these defined components of the ETH metric characterization, rather than denoting direct matrix elements of physical observables.
@@ III.1 Trade-off between diagonal and off-diagonal elements
For any ensemble $\rho_{B}$ with strictly positive reduced matrix $\rho_{B_{1}}=\text{Tr}_{\bar{B}_{1}}\rho_{B}$, the diagonal element $V_{\text{dg}}^{i}:=V(\rho_{B},O^{B_{1}}_{ii})$ and the off-diagonal element $V_{\text{off}}^{ij}:=V(\rho_{B},O^{B_{1}}_{ij;i\neq j})$ in $d^{(3)}$ are not independent. According to Appendix A, they are related by the following equation
$V_{\text{dg}}^{i}+\sum_{j\neq i}V_{\text{off}}^{ij}=\text{Tr}(\rho_{B_{1}}^{-1})-1,$ (22)
where the index $j$ runs over all states in the Hilbert space excluding $i$. The summation contains $D_{N}-1$ terms, whose number scales exponentially with the system size $N$. However, the right-hand side of (22) depends solely on the local state $\rho_{B_{1}}$, and it generally grows exponentially with the size of subsystem $N_{B_{1}}$, while remaining relatively constant with respect to $N$. This causes the off-diagonal elements to be exponentially suppressed. Considering the average of these off-diagonal elements
$V_{\text{off}}^{i}:=\frac{1}{D_{N}-1}\sum_{j\neq i}V_{\text{off}}^{ij},$ (23)
which will decrease exponentially with $N$. However, this exponential decrease in the average of the off-diagonal elements does not imply that the maximum values of the off-diagonal elements and the diagonal element will also decrease exponentially. Although the diagonal elements are not directly suppressed, Eq. 22 provides a trade-off relation between the diagonal element and the off-diagonal elements. With $V_{\text{dg}}^{i}/V_{\text{off}}^{i}=:f^{i}(N)$, we can rewrite Eq. 23 as
@@
On one hand, if the ratio between diagonal and average off-diagonal elements $f^{i}(N)$ does not increase exponentially, or grows exponentially at a rate much slower than $D_{N}$, then in the large $N$ limit, the diagonal elements will also be exponentially small. On the other hand, if the diagonal element is not decreasing exponentially, $f^{i}(N)$ must be growing exponentially. Thus, the exponential growth of $f^{i}(N)$ is a necessary condition for the diagonal element to avoid exponential decay.
@@
$f^{i}(N)\leq\frac{D_{N}-1}{\sum_{j\neq i}\lVert(\sigma^{i}_{B_{1}}-\rho_{B_{1}})(\sigma^{ij}_{B_{1}})^{-1}\rVert^{-2}_{\infty}},$ (26)
@@
where $|\sigma_{B_{1}}^{ij}=\text{Tr}_{\bar{B}_{1}}\ket{E_{i}}\bra{E_{j}}$ |. The right-hand side of inequality (26) may provide a more straightforward framework to discuss the effects of interactions and the choice of the $\rho_{B_{1}}$ on $f^{i}(N)$.
@@
Figure 1: Diagonal and off-diagonal elements versus eigenstate index. (a) Spin chain under PBC. (b) Spin chain under OBC. In the plot, $\text{Var}_{\text{dg}}:=V_{\text{dg}}^{i}$, $\text{Var}_{\text{off}}:=\sum_{j\neq i}V_{\text{off}}^{ij}$, $\text{SN}_{\text{dg}}:=\lVert\sigma_{B_{1}}^{i}-\rho_{B_{1}}\rVert_{1}^{2}$ and $\text{SN}_{\text{off}}:=\sum_{j\neq i}\lVert\sigma_{B_{1}}^{ij}\rVert_{1}^{2}$. The variance measures (Var) systematically produces larger values than the trace-norm measures (SN), validating Section II.1 and the hierarchy in Table 1. The measure $\text{Var}_{\text{dg}}$ closely aligns with $\text{SN}_{\text{dg}}$, making the data points for $\text{Var}_{\text{dg}}$ overlap with those for $\text{SN}_{\text{dg}}$. Eigenstates are sorted by energy (low to high). Parameters: $g=1.05$, $h=0.1$, $N=12$, $N_{B_{1}}$=1, $\rho_{B}=\sum_{i}\Pi^{i}_{B}/D_{N}$.
@@
The index $i$ in Eq. 22 can be chosen arbitrarily, but the right side of the formula is independent of eigenstate index $i$. When the measure $V_{\text{dg}}$ of a given eigenstate is large, the average measure of its off-diagonal elements $V_{\text{off}}^{i}$ must be small, and vice versa. Fig. 1 illustrates this trade-off relation, where the energy eigenstates are numerically calculated based on a specific non-integrable system. We will introduce the model used in detail in Section IV. Since the measure $d^{(2)}$ and $d^{(3)}$ are related by Section II.1, the trade-off between the diagonal and off-diagonal elements is also reflected in $d^{(2)}$, although it is not an exact equation relationship like Eq. 22. From the Fig. 1, we observe that the diagonal element is generally much smaller than the sum of the corresponding off-diagonal elements. But this does not mean that $f^{i}$ is generally much smaller than 111. On the contrary, since there are $D_{N}-1$ off-diagonal elements for each diagonal element, $f^{i}$ is typically much larger than 111, which will be shown in Fig. 2.
@@
The formula (22) holds for any $\rho_{B}$ as long as $\rho_{B_{1}}$ is strictly positive. This includes not only the canonical ensemble $\rho^{\text{c}}(\beta):=e^{-\beta H}/Z(\beta)$ and microcanonical ensembles, but also pure states such as $\Pi^{i}_{B}$. In the case of pure states, the diagonal element in Eq. 22 will be zero, leading to
$\sum_{j\neq i}V(\Pi^{i}_{B},O^{B_{1}}_{ij})=\text{Tr}(\sigma_{B_{1}}^{-1})-1,$ (27)
@@
where the sum of all corresponding off-diagonal elements is directly determined by the local state $\sigma_{B_{1}}=\text{Tr}_{\bar{B}_{1}}\Pi^{i}_{B}$.
The trade-off relationship is not exclusive to the measure $d^{(3)}$; Other measures can also yield similar relationships. For instance, for measure $d^{(1)}$, it can be shown that
$\sum_{j}[d^{(1)}(\Pi^{ij}_{B},\rho_{B};A^{B_{1}})]^{2}=\sum_{j}\lvert\bra{E_{j}}A^{B_{1}}-\braket{A^{B_{1}}}\ket{E_{i}}\rvert^{2}$
$=\text{Tr}[\Pi^{i}_{B}(A^{B_{1}}-\braket{A^{B_{1}}})(A^{B_{1}}-\braket{A^{B_{1}}})^{\dagger}].$ (28)
Compared with Eq. 22, the right-hand side of Section III.1 depends on the state index $i$. However, since $(A^{B_{1}}-\braket{A^{B_{1}}})(A^{B_{1}}-\braket{A^{B_{1}}})^{\dagger}$ is a local operator on $B_{1}$, when ETH holds, the right-hand side of the equation will approximate the variance $V(\rho_{B},A^{B_{1}})$ according to Eqs. 2 and 8.
Another key difference lies in whether the measure is observable-independent. The right-hand side of Section III.1 explicitly depends on the observable $A^{B_{1}}$, contrasting with the observable-independent nature of the measure in Eq. 22. For the observable-independent measure $d^{(2)}$, its form is not easily transformed, and thus we have not yet identified a direct trade-off relationship. However, as shown in the numerical calculations for a specific model in Fig. 1, a similar relationship does emerge, though, like in Section III.1, it is not independent of the state index $i$. For the measure $d^{\prime(2)}$, the following trade-off holds:
$\sum_{j}[d^{\prime(2)}(\sigma_{B_{1}}^{ij},\rho_{B_{1}})]^{2}$
$=\sum_{j}\text{Tr}(\sigma^{ij}_{B_{1}}\sigma^{ij{\dagger}}_{B_{1}})+\text{Tr}(\rho_{B_{1}}^{2})-2\text{Tr}(\rho_{B_{1}}\sigma_{B_{1}}^{i})$
$=D_{B_{1}}+\text{Tr}(\rho_{B_{1}}^{2})-2\text{Tr}(\rho_{B_{1}}\sigma_{B_{1}}^{i}),$ (29)
where the final equation follows from substituting $\rho_{B_{1}}^{-1}=I_{B_{1}}$ in Appendix A. Here, $D_{B_{1}}$ denotes the Hilbert space dimension of subsystem $B_{1}$. The right-hand side of Section III.1 also depends on the state index $i$.
@@ III.2 Weak ETH
We examine the use of measure $d^{(3)}$ to study the weak ETH with eigenstate typicality by considering both “diagonal” and “off-diagonal” probabilistic typicalities. The “diagonal” probabilistic typicality is given by
@@
while the “off-diagonal” probabilistic typicality is
@@
Similar to the discussion in the previous section, the probability average of off-diagonal elements $\braket{V_{\text{off}}}$ should decrease exponentially with $N$. According to Eq. 24, the “diagonal” probabilistic typicality
$\braket{V_{\text{dg}}}=[\text{Tr}(\rho_{B_{1}}^{-1})-1]\times\sum_{i}p_{i}\frac{f^{i}(N)}{f^{i}(N)+D_{N}-1},$ (33)
and its decay behavior with increasing system size depends on the specific model.
@@ III.3 The average observable
If we consider the distinguishability measure $V(\rho_{B},\overline{O}^{B_{1}}_{ij})$ given by the average observable $\overline{O}^{B_{1}}_{ij}=\frac{1}{\mathcal{C}}\sum_{k}\mathcal{J}_{\rho_{B_{1}}}^{-1/2}(\sigma^{ij,k}_{B_{1}})$, then according to Appendix A, the corresponding diagonal and off-diagonal elements exhibit the following trade-off relationship
$V(\rho_{B},\overline{O}^{B_{1}}_{ii})+\sum_{j\neq i}V(\rho_{B},\overline{O}^{B_{1}}_{ij})=\frac{1}{\mathcal{C}}(\text{Tr}(\rho_{B_{1}}^{-1})-1)$
$+\frac{1}{\mathcal{C}^{2}}\sum_{\beta,\alpha,l,k\neq l}p^{\prime}_{\alpha}\text{Tr}[O^{B_{k},{\dagger}}_{\alpha\beta,1}\otimes O^{B_{l}}_{\alpha\beta,1}(\sigma^{i}_{B_{k}B_{l}}-\sigma^{i}_{B_{k}}\otimes\rho^{1}_{B_{l}})],$ (34)
where $\rho^{1}_{B_{l}}$ is the translational copy of $\rho_{B_{1}}$. Although formula (III.3) is valid for any $\rho_{B}$ with strictly positive $\rho_{B_{1}}$, it connects to Section II.4 specifically when using $\overline{\rho}_{B}$. Thus, all quantities related to $\rho_{B}$ should be replaced with $\overline{\rho}_{B}$, which is translation invariant. In such cases $\overline{\rho}^{1}_{B_{l}}=\overline{\rho}_{B_{l}}$. The local term in Section III.3 becomes $\text{Tr}(\overline{\rho}_{B_{1}}^{-1})-1$, which is related to the local state $\overline{\rho}_{B_{1}}$ and includes an additional factor of $1/\mathcal{C}$. The local term decreases algebraically as $\mathcal{C}$ increases. The second term on the right side of Section III.3 is related to the spatial correlation of $\sigma^{i}$ and the difference between $\sigma^{i}_{B_{l}}$ and $\rho^{1}_{B_{l}}$ (or $\overline{\rho}_{B_{l}}$ for translation invariant system).
@@
$\braket{V_{\text{dg}}}=\sum_{i}p_{i}V(\rho_{B},\overline{O}^{B_{1}}_{ii}).$ (35)
@@
$\braket{V_{\text{off}}}=\frac{1}{D_{N}-1}\sum_{i,j\neq i}p_{i}V(\rho_{B},\overline{O}^{B_{1}}_{ij}).$ (36)
@@
$V^{\text{Avg}}_{\text{Cor}}:=\frac{1}{\mathcal{C}^{2}}\sum_{\beta,\alpha,l,k\neq l}p^{\prime}_{\alpha}\text{Tr}[O^{B_{k},{\dagger}}_{\alpha\beta}\otimes O^{B_{l}}_{\alpha\beta}(\rho_{B_{k}B_{l}}-\rho_{B_{k}}\otimes\rho_{B_{l}})],$ (38)
which is determined by the spatial correlation of $\rho_{B}$.
As previously mentioned, when using Section II.4, all quantities in formula (37) related to $\rho_{B}$ must be replaced with $\overline{\rho}_{B}$, except for $\rho_{B_{k}B_{l}}$ and $\rho_{B_{k}}$ in Eq. 38. If $\rho_{B}$ is translation invariant, then $\rho_{B}=\overline{\rho}_{B}$, and the second term on the right side of Eq. 37 is solely related to the correlation of $\overline{\rho}_{B}$. If the spatial correlation of $\overline{\rho}_{B}$ decays exponentially or algebraically fast enough, the right side of Eq. 37 will exhibit algebraic decay of the form $O(N^{-\epsilon})$ [18]. Compared to Eq. 32, the right side of Eq. 37 depends not only on the local state but also on the rate of correlation decay. Both terms generally decrease algebraically with increasing $N$, consistent with the intuition that spatial averaging smooths and reduces differences (distinguishability). Although this algebraic decay alone does not guarantee strong ETH, it supports the validity of weak ETH with algebraic decay, even for integrable systems. This algebraic decay aligns with the findings of our previous article [18], though we did not previously account for the $D_{N}-1$ off-diagonal elements in Eq. 37, which cause the average off-diagonal elements to be exponentially suppressed and further restrict their values. Consequently, as the right side of Eq. 37 decreases with $N$, the exponential decrease of $\braket{V_{\text{off}}}$ remains applicable. Additionally, due to the smoothing effect of the average observable, the values of $\braket{V_{\text{dg}}}$ and $\braket{V_{\text{off}}}$ here will be smaller than those defined in the Section III.2.